1. n=8, p =0.59 to find the following Binomial probability P(0) = NOTE: Round your answer...

1.

n=8, p =0.59

to find the following Binomial probability

P(0) =

NOTE: Round your answer to THREE decimal places.

2. The community church is having a fundraising to send its kids to a field trip. The church sells 401 tickets for a raffle. The price of the tickets are $3. If the prize of this raffle is 244, what is the expected value of this raffle?

NOTE: Round your answer to TWO decimal places.
Answer #1

#1.
Here, n = 8, p = 0.59, (1 - p) = 0.41 and x = 0
As per binomial distribution formula P(X = x) = nCx * p^x * (1 - p)^(n - x)

We need to calculate P(X = 0)
P(X = 0) = 8C0 * 0.59^0 * 0.41^8
P(X = 0) = 0.001


#2.

Probability that a person wins the price is 1/401

Expected value = 244*(1/401) - 3*(400/401) = -2.38